Let f: R^2 to R be a function defined by

f(x,y,z) = { 1/(x-1) + 1/(y-1)+1/(z-1);

x≠1,y≠1,z≠1

{ 0 ; Otherwise

Calculate fxz(1,1,1)


1
Expert's answer
2021-06-17T12:26:50-0400

Given f(x,y,z) = "\frac{1}{x-1} + \frac{1}{y-1} +\frac{1}{z-1}"  for "x\neq 1, y\neq 1, z\neq 1" and otherwise f(x,y,z) = 0

fx(x,y,z) = "\frac{\delta}{\delta x}(\frac{1}{x-1} + \frac{1}{y-1} +\frac{1}{z-1} )"


fx(x,y,z) = "\frac{-1}{(x-1)^2}"


fxz(x,y,z) = "\frac{\delta}{\delta z}" fx(x,y,z) = "\frac{\delta}{\delta z}" "(\frac{-1}{(x-1)^2})" = 0

therefore,

"\implies" fxz(x,y,z) = 0 for all (x,y,z) "\in" R3

hence,

fxz(1,1,1) = 0